The average cost of a certain type of grass seed is Birr 4.00. The standard deviation is 

0.1 Birr. Find the minimum percentage of data values, which fall in the 3.82 to 4.18

range.


1
Expert's answer
2022-06-13T16:30:38-0400

"P(3.82<X<4.18)=P(\frac{3.82-4}{0.1}<Z<\frac{4.18-4}{0.1})=P(-1.8<Z<1.8)="

"=P(Z<1.8)-P(Z<-1.8)=0.9282."